# 3D-Properties: Identifying Challenges in DPO and Charting a Path Forward

### v1

@@ 3.2.1 Toy Model Setup
We construct discrete spaces consisting of 4 prompts and 10 responses. The policy $\pi_{\theta}$ is implemented as a three-layer MLP that processes a one-hot vector to produce a categorical distribution over the responses. The first 4 dimensions of the response space correspond to the chosen responses, dimensions 5 through 8 to the rejected responses, and the final 2 dimensions to unseen responses not present in the preference dataset. We configure the optimal response to correspond directly with its respective prompt; for example, response 1 is optimal for prompt 1. This setup is illustrated in Figure 1.
Step 1. We simulate the Pretraining/SFT process by manually specifying the output probabilities and using them as labels to train the model.
Step 2. We construct the preference dataset and simulate the DPO process. The dataset is constructed using optimally chosen responses and randomly selected rejected responses to form paired data.
@@ 3.2.2 Results
Initially, we set the likelihood of each chosen and rejected response at $0.12$, allowing us to treat both as on-policy. The results are shown in Figure 2. During the early phases of DPO, the likelihood of chosen responses increases, while that of rejected responses decreases. However, as training progresses and $\pi_{\theta}(a^{-}|x)$ approaches zero (illustrated in the third figure), the likelihood of chosen responses begins to diminish. Concurrently, in the second figure, we have plotted the changes in gradient, which shows that the absolute value of $\partial\ell^{\text{DPO}}/\partial\pi^{-}$ increases sharply. In the degradation phase, with both chosen and rejected responses decreasing in likelihood, the probability is reassigned to unseen responses. These findings are consistent with our previous theoretical analysis.
We now understand that the rate at which $\pi_{\theta}(a^{-}|x)$ declines is a critical factor in determining the severity of the 3D-properties’ impact. This leads us to propose the following:
@@
Furthermore, this insight helps explain why on-policy DPO is generally more effective (Fact.3). Compared to off-policy DPO, on-policy DPO starts with a higher likelihood of rejected responses, thereby prolonging the time it takes for their likelihood to diminish. On the other hand, the higher initial probability within the training set also minimizes the dispersion effect on unseen tokens.
Figure 3: From left to right, the figures present the initial state and the changing process of the likelihood of each response for Scenario 1 to Scenario 4. Scenario 1: chosen on-policy and rejected on-policy. Scenario 2: chosen off-policy and rejected on-policy. Scenario 3: chosen on-policy and rejected off-policy. Scenario 4: chosen off-policy and rejected off-policy.
To empirically validate these differences between on-policy and off-policy DPO, we conducted experiments using our toy model. We configure four scenarios by adjusting the initial distribution of the outputs in Step 1. The initial state and subsequent changes in the likelihood of each response are depicted in Figure 3. Notably, the 3D-properties in Scenario 1, where both chosen responses and rejected responses are on-policy, are relatively mild, as evidenced by both the high peak probability of the optimal response (approximately $0.6$) and the minimal dispersion effect on unseen responses. Further detailed results and analyses of these four scenarios are provided in Appendix D.
3.3 Regularization Techniques
Inspired by Proposition 1, we introduce two straightforward regularization techniques. The first technique employs variable values of $\beta$ to control the rate at which the likelihood of rejected responses declines. The second technique involves augmenting the DPO loss with an SFT loss, a strategy that has been shown to significantly enhance the stability of DPO in previous studies, such as those documented in [15, 32]. These regularization methods have shown promising results in our toy model. Detailed discussions and experimental outcomes are presented in Appendix B.2.
4 Non-existence of 3D-properties in RLHF-PPO
In this section, we show that 3D-properties do not manifest in the commonly used alignment methods RLHF-PPO, which may account for why DPO methods only achieve sub-optimal performance in comparison.
In the RLHF-PPO algorithm described in [21], the process is twofold: first, training a reward model (RM), followed by optimization using a reinforcement learning algorithm such as PPO. The loss function for the RM is defined as:
$\ell^{\text{RM}}(\theta)=-\sum_{(x,a^{+},a^{-})}\log\sigma\left(r(a^{+}|x)-r(a^{-}|x)\right).$
For a given $(x,a^{+},a^{-})$, let
$r^{+}:=r(a^{+}|x),\quad r^{-}:=r(a^{-}|x),$
the gradients with respect to $r^{+}$ and $r^{-}$ are:
@@ 4 Non-existence of 3D-properties in RLHF-PPO
This indicates that the gradients for the chosen and rejected responses are balanced and do not exhibit 3D-properties. In Section 5.2.1, we will further discuss the relationship between DPO and RLHF-PPO in real LLMs.
@@ 5 Experiments
In this section, we transition from theoretical analyses and toy model explorations to conducting real-world experiments with LLMs. These experiments are designed to validate the theoretical insights detailed in Section 3. The experimental setup is outlined in Section 5.1, with the main experimental results presented in Section 5.2 and a comparative analysis in Section 5.3.
@@ 5.1 Experimental Setup
Datasets. The alignment phase is commonly conducted to improve the model’s capabilities in mathematical reasoning and instruction following. Accordingly, we first validate our theory in the field of mathematics using the MATH dataset [14] as our primary experimental dataset [^1] Only the questions from the dataset were used to generate the preference dataset; details are provided in Section 5.2.. Additionally, to evaluate the model’s out-of-distribution (OOD) generalization capabilities, we employed the SuperCLUE-Math benchmark [30]. To assess instruction-following abilities, we also created two in-house preference datasets focused on poem and slogan generation.
In practical applications, the commonly used datasets for the alignment stage cover various domains, and the sources of data can also be diverse. Thus, we further use a self-built general dataset consisting of approximately 400,000 preference samples, covering a diverse range of domains and the source of the data is also various, ranging from data generated by the model itself, data generated by GPT, and data from open-source datasets like HH-rlhf [6]. We designate this self-built dataset as "COMMON". The size of these datasets is detailed in Table 7 in the Appendix.
@@
The LLMs of concern. We focus on Baichuan2-13B and Baichuan2-33B, an advanced bilingual (Chinese and English) LLM series. The 13B model is openly available [33], and the 33B model extends the 7B architecture with increased parameters.
5.2 The Effect of the Distribution of the Training Data.
Building on the theoretical insights presented in Section 3, we hypothesize that the performance of the vanilla DPO algorithm is influenced by the distribution characteristics of the training data, specifically whether they are on-policy or off-policy. For a detailed discussion on on-policy versus off-policy training, refer to Section A.1.
@@
To generate on-policy responses, we used the model to be aligned to produce 8 candidates for each prompt in the MATH training set. Then, using GPT-4 [1], we selected the best and the worst responses from 8 candidates to form a preference pair. After applying filtering operations to exclude cases where all responses were uniformly good or bad, we compiled the MATH dataset, which comprises 5826 training triplets ${x,a^{+},a^{-}}$. We randomly select 2,000 samples from the original test set as the test set for MATH. For off-policy chosen responses, we utilized the original solutions provided in the dataset. For the off-policy rejected responses, we employed responses generated by the Qwen1.5-7B model [5] for each prompt.
To validate the hypothesis, we implemented the settings described in the previous 4 scenarios shown in Figure 3. We used DPO to train the model under these scenarios respectively. After training, we test the performance of the model with the help of GPT-4. GPT-4 will assign scores ranging from 1-5 to each response, with responses scoring above 4 considered correct answers. The scoring criteria are detailed in LABEL:tab:score_criteria. The average performance results on the MATH and SuperCLUE-Math datasets are reported in Table 2. The specific experimental results for the two datasets are available in Tables 8 and 9. Additionally, we report the log probabilities before and after training in Table 1.
@@
Among all 4 scenarios, Scenario 1 features both chosen and rejected responses as on-policy. This configuration ensures a more stable DPO training process, resulting in the best testing performance, as is shown in Table 2. According to Proposition 1, the critical factor determining the impact of 3D-properties is the decline rate of the likelihood of the rejected responses, $\log\pi(a^{-})$. As is shown in Table 1, in comparison to other scenarios, Scenario 1 shows the slowest decline in likelihood, effectively mitigating the adverse effects of 3D-properties. This explains the superior performance of on-policy DPO in our tests.
@@
We also plot the gradients during the training process of the DPO algorithm in scenario 1, which is shown in Figure 10 in the Appendix. This visualization corroborates the analysis presented in Section 3, demonstrating that the gradients for rejected responses increase more rapidly during training. Such an excessive decrease in the likelihood of generating rejected responses can ultimately contribute to model degradation. Additional analysis of the other scenarios is provided in Appendix D.
Table 1: The impact of training data being on-policy on the results. $\log\pi(a^{+})$ and $\log\pi(a^{-})$ denote the average log probability of each token of chosen response $a^{+}$ and rejected response $a^{-}$.
Baichuan2-13B Baichuan2-33B
$\log\pi(a^{+})$ $\log\pi(a^{-})$ $\log\pi(a^{+})$ $\log\pi(a^{-})$
basemodel -0.9181 -0.9393 -0.3603 -0.3634
DPO in Scenario 1 -0.9681 -0.9982 -0.3629 -0.3670
basemodel -1.6776 -0.9238 -1.4314 -0.3525
DPO in Scenario 2 -1.6265 -1.2293 -1.2734 -0.4254
basemodel -0.9461 -1.7110 -0.3460 -1.1204
DPO in Scenario 3 -0.8786 -1.8848 -0.3439 -1.4333
basemodel -1.7468 -1.7110 -1.2838 -1.1451
DPO in Scenario 4 -1.6617 -1.8265 -1.2273 -1.2234
Table 2: The performance tested on MATH test set and SuperCLUE-Math.
@@
Table 3: The test result on self-built dataset Poem and Slogan. All the metrics are such that higher values are better.
@@
5.2.1 Suboptimality of DPO Algorithm Compared to RLHF-PPO Algorithm
@@
In addition to mathematics datasets, we also tested the DPO algorithm and the commonly used RLHF-PPO algorithm on two self-built datasets regarding poem creation and slogan creation, respectively. These datasets have explicit scoring standards. For example, in the scenario of creating poems, the model needs to generate responses in different text and tone formats based on the given prompt. We set the metrics of a poem to the following five aspects: Row Number, Words per Row, Rhythm, Tone Pattern and Title, and set the metrics of a slogan to the following two aspects: Word Count and Content. For a detailed explanation of each metric, see Appendix C for detailed information. We use Baichuan2-33B to conduct the experiments and the results are shown in Table 3. As can be seen, the performance of DPO on both two datasets is worse than that of RLHF-PPO.
Following the regularization methods mentioned in Section 3.3, we fixed $\beta_{1}$ and gradually decreased $\beta_{2}$. The results are shown in Figure 9 in the Appendix. As $\beta_{2}$ decreases, performance in poem generation initially improves and subsequently declines. The initial improvement is intuitive and aligns with our expectations, as lowering $\beta_{2}$ slows the decrease of $\pi_{\theta}(a^{-}|x)$. However, we conjecture that an excessively low $\beta_{2}$ may cause DPO to perform like SFT on the chosen responses, thereby reducing its generalization capabilities. Additionally, we observe that other DPO variants like IPO and SLiC, despite having less severe 3D-properties effects, perform worse than DPO and PPO, which echoes Fact 2. One hypothesis is that their solution forms diverge significantly from the Bradley-Terry model, leading to a loss of generalization in preference learning. Both aspects demand further investigation. We provide a preliminary discussion in Appendix D and take them as parts of our future research.
5.3 Relative Instability of DPO Training Compared to Reward Model Training
(a) Accuracy of reward model and DPO on HH-rlhf eval set over the training process.
(b) Accuracy of reward model and DPO on MATH eval set over the training process.
Figure 4: Comparison between DPO and reward model training.
Following Section 4, in this section, we conduct a direct comparison between DPO training and reward model (RM) training. Given that the objective function for DPO is derived from that used to train an RM, we simultaneously train an RM and conduct DPO training, allowing us to explicitly demonstrate the relative instability of DPO training.
We employ the Baichuan2-33B model for both DPO training and the initialization of the RM. The training preference dataset is our self-constructed COMMON dataset, and we test the two models on the widely used HH-rlhf dataset [6] and the MATH dataset [14]. We adopt accuracy as our metric, defined as the proportion of instances in which the model correctly identifies the chosen response as superior to the rejected one.
As is shown in Figure 4, the training process of the reward model is more stable, while the process of the DPO training fluctuates significantly. This observation aligns with the theoretical findings presented in Section 4, which state that 3D-properties do not exist in reward-based alignment methods. Furthermore, results from Figure 11 and Figure 12 show that the model trained using DPO is more likely to overfit the training data, reflecting a more aggressive optimization strategy.
6 Concluding Remarks and Limitations

### v2

@@ 3.2.1 Toy Model Setup
The diagram for the toy model is illustrated in Figure 1. We construct a discrete space consisting of 4 prompts and 10 responses. The policy $\pi_{\theta}$, which simulates a simplified version of an LLM, is implemented as a three-layer MLP that processes a one-hot vector and outputs a categorical distribution over the responses. The response space is organized such that the first 4 dimensions correspond to chosen responses, dimensions 5 through 8 represent rejected responses, and the final 2 dimensions correspond to unseen responses not present in the preference dataset.
In this setup, each prompt has an optimal response (e.g., response 1 is optimal for prompt 1, as shown in the upper left figure). When constructing the preference dataset for DPO training, we adopt a mini-batch sampling strategy to mimic real-world annotation processes. Specifically, assuming an ideal annotator, each input prompt is perfectly matched with its optimal response—corresponding to the diagonal elements of the matrix, as illustrated in the upper right figure in Figure 1. For each mini-batch, we then randomly select one other response within the batch to create preference data pairs. This approach ensures that gradient updates are computed from diverse mini-batch samples.
To simulate the Pretraining and SFT process, we manually assign output probabilities and use them as labels to train $\pi_{\theta}$. Initially, as shown in the lower left figure, we set the likelihood of both chosen and rejected responses at $0.12$, treating both as on-policy. The constructed preference dataset is then used for DPO training, with the output after 500 epochs shown in the lower right. The code is provided in the supplementary material.
@@ 3.2.2 Results
Figure 2 illustrates the dynamic optimization process during DPO training. In the first figure, the likelihood of chosen responses (blue and yellow curves) increases, while the likelihood of rejected responses (green and red curves) decreases in the early phases of training. However, as training progresses, the likelihood of chosen responses begins to decline in the longer run. During this degradation phase, as both chosen and rejected response likelihoods decrease, the probability is redistributed to unseen responses (purple curve).
The second and third figures reveal the underlying causes of this shift: as $\pi_{\theta}(a^{-}|x)$ approaches zero, the absolute value of $\partial\ell^{\text{DPO}}/\partial\pi^{-}$ increases sharply compared to $\partial\ell^{\text{DPO}}/\partial\pi^{+}$. The absolute value of $\partial\ell^{\text{DPO}}/\partial\pi^{+}$ becomes progressively smaller, weakening its influence on the optimization direction. These results align with the earlier theoretical analysis.
Figure 3: From left to right, the figures show the initial state and the likelihood dynamics for chosen/rejected/unseen responses in Scenarios 1 to 4, similar to the left diagram in Figure 2: (1) both chosen and rejected responses are on-policy, (2) chosen off-policy and rejected on-policy, (3) chosen on-policy and rejected off-policy, and (4) both off-policy.
Moreover, this insight provides an explanation for the observed superiority of on-policy DPO (Observation 3). In contrast to off-policy DPO, on-policy DPO begins with a higher likelihood for rejected responses, thereby extending the duration before their likelihood significantly diminishes. To further validate the differing impacts of on-policy and off-policy DPO, we configure four scenarios by adjusting the initial distribution of outputs to simulate these conditions. A higher initialized likelihood (0.12) simulates responses sampled in an on-policy manner, while a lower one (0.02) simulates responses sampled off-policy. The initial state and the subsequent changes in the likelihood of each response are illustrated in Figure 3. Notably, in Scenario 1, where both chosen and rejected responses are on-policy, the 3D-properties are relatively mild, as shown by the high peak probability of the optimal response (approximately $0.6$) and the minimal dispersion effect on unseen responses. Additional detailed results and analyses for all four scenarios can be found in Appendix D.
The intention of the toy model and its connection to real LLMs. The toy model serves as an abstract simulation that amplifies the effect of 3D-properties, which are less pronounced and harder to visualize in real-world experiments. While the toy model differs from real LLM training in several ways—such as sampling frequency—its design offers useful insights. In real-world settings, DPO is typically trained over one epoch, with each data point used only a few times. In contrast, in the toy model, the same data points are sampled repeatedly. Conceptually, this is similar to treating each input/output as a token rather than a complete prompt/response, where each token may be sampled multiple times during real-world training. Since both the chosen and rejected responses are generated from the same prompt, they often share common tokens. As a result, the decrease in the likelihood of rejected responses can impact the likelihood of chosen responses, leading to a corresponding decline in their likelihood.
3.3 Regularization Techniques
It becomes evident that the rate at which $\pi_{\theta}(a^{-}|x)$ declines is crucial in determining the severity of the 3D-properties’ impact. This observation leads to the following proposition:
@@ 3.3 Regularization Techniques
Inspired by Proposition 1, we introduce two straightforward regularization techniques. The first technique employs adaptive values of $\beta$ to control the rate at which the likelihood of rejected responses declines, referred as Flex-DPO. The second technique involves augmenting the DPO loss with an SFT loss, a strategy that has been shown to significantly enhance the stability of DPO in previous studies (Hou et al., 2024; Xu et al., 2024b). These regularization methods have shown promising results with our toy model and will be further validated in real LLMs in the following section. The theoretical analysis is similar to that of vanilla DPO thus deferred to Appendix B.2.
3.4 Inherent Absence of 3D-properties in RM-based alignment
In this section, we show that 3D-properties do not manifest in RM-based alignment methods, which may account for why DPO methods only achieve sub-optimal performance. Since DPO is closely related to RM training, and the Best-of-N performance of the RM can partially reflect the ultimate performance of the policy model (Gui et al., 2024), we focus on analyzing the RM’s objective. For a given $(x,a^{+},a^{-})$, let $r^{+}:=r(a^{+}|x)$ and $r^{-}:=r(a^{-}|x)$, the gradients with respect to $r^{+}$ and $r^{-}$ are:
@@ 3.4 Inherent Absence of 3D-properties in RM-based alignment
This indicates that the gradients for the chosen and rejected responses are balanced and do not exhibit 3D-properties. In Section 4.5, we will further discuss the relationship between DPO and RM-based alignment in real LLMs.
@@ 4 Experiments
In this section, we transition from theoretical analyses and toy model simulations to real-world experiments with LLMs to further validate our theoretical insights. We verify the existence of 3D-properties, the superiority of on-policy DPO over off-policy DPO, the superiority of RM over DPO, and the effectiveness of the proposed regularization technique.
@@ 4.1 Experimental Setup
Datasets. We chose mathematical reasoning and instruction following as our primary benchmarks because these tasks are easily quantifiable, providing clear metrics for evaluating model performance. For mathematical reasoning, we used MATH (Hendrycks et al., 2021) as the main dataset for both training and testing [^1] Only the prompts from the dataset were used to generate the preference dataset; further details are provided in Section 4.2.. To assess the model’s out-of-distribution (OOD) generalization capabilities, we selected SuperCLUE-Math (Xu et al., 2020), another dataset which was used exclusively for testing. Additionally, we included two in-house datasets focused on poem and slogan generation to evaluate the model’s ability to handle creative tasks with strict structural and linguistic constraints. The poem dataset, for instance, which has rigid format and rhyme requirements, making it a good test for evaluating the model’s ability to follow complex instructions. Further details and descriptions of the datasets used are provided in Appendix C.
It is widely accepted in industry that preference datasets for model alignment should cover a broad range of domains. Following this consensus, we further utilized a general dataset consisting of approximately 400,000 preference samples across diverse domains. These prompts were sourced from HH-rlhf (Bai et al., 2022) and UltraFeedBack (Cui et al., 2024). A detailed breakdown of the dataset sizes is provided in Table 6 in Appendix C.
@@
The LLMs of concern. We focus on Baichuan2-13B and Baichuan2-33B, an advanced bilingual (Chinese and English) LLM series. The 13B model is openly available (Yang et al., 2023), and the 33B model extends the 13B architecture with increased parameters.
4.2 Effect of Training Data Distribution: on-policy vs. off-policy
Building on the theoretical insights in Section 3, we hypothesize that the performance of vanilla DPO is significantly influenced by the distribution gap between the training dataset and outputs of the policy model, specifically whether the algorithm is on-policy or off-policy. Off-policy DPO uses an external preference dataset, while on-policy DPO samples preferences directly from the policy model. On-policy DPO enjoys a smaller distribution gap compared with off-poliy DPO.
@@
To conduct on-policy DPO, we used the policy model to produce 8 candidates for each prompt in the train set of MATH. The best and worst responses were selected by GPT-4 (Achiam et al., 2023) to form a preference pair, with the standard solutions given as the reference context. After filtering out uniformly good or bad responses, we compiled the MATH dataset, which contains 5,826 pairs $\{x,a^{+},a^{-}\}$. We randomly selected 2,000 samples from the original test set to serve as the test set for MATH. For off-policy DPO, we used the original solutions from the dataset as the chosen responses, and generated the rejected responses using Qwen1.5-7B (Bai et al., 2023), a relatively earlier model with limited capabilities.
To validate the hypothesis that the presence of off-policy data weakens performance, we implemented the four scenarios consistent with the toy model (Figure 3). We evaluated the policy model using GPT-4, which assigned scores ranging from 1 to 5 based on the accuracy of both the final answer and the problem-solving process, with also the standard solutions given as the reference context. The scoring criteria are detailed in LABEL:tab:score_criteria. The average performance on the MATH and SuperCLUE-Math datasets is reported in Table 1, with specific results in Table 8. Among the four scenarios, Scenario 1—where both chosen and rejected responses are on-policy—ensured a more stable DPO training process and delivered the best performance.
@@
Additionally, we report the log probabilities before and after training in Table 7 in Appendix D.2. According to Proposition 1, the key factor affecting the impact of 3D-properties is the decline rate of rejected responses’ likelihood, $\log\pi(a^{-})$. Scenario 1 shows the slowest decline in likelihood compared to the other scenarios, effectively mitigating the adverse effects of 3D-properties, which explains the superior performance of on-policy DPO in our tests.
@@
We also plot the gradients during the DPO training process for Scenario 1, as shown in Figure 6 in the Appendix D.2. This visualization supports the analysis in Section 3, demonstrating that the gradients for rejected responses increase more rapidly during training. This excessive decline in the likelihood of generating rejected responses can ultimately lead to model degradation.
In addition to pure on-policy and off-policy DPO, Scenario 2, where the chosen response is off-policy and the rejected response is on-policy, is also prevalent in industry. For instance, in math problems, researchers often treat the correct dataset solution as the chosen response and the LLM-generated incorrect answer as the rejected one, which we demonstrate to be detrimental. Scenario 3 is a mirror experiment for Scenario 2. These experiments confirm that incorporating off-policy data into the preference training set degrades DPO performance.
Table 1: Results tested on MATH and SuperCLUE-Math. Scenario 1, with both chosen and rejected responses sampled on-policy, shows the best performance.
@@
Table 2: Test results on the self-built Poem and Slogan datasets. All metrics are evaluated such that higher values indicate better performance.
@@
4.3 Experimental Validation of Regularization Techniques
Figure 4: Performance on poem generation, $\beta^{-}$ varying with $\beta^{+}=0.1$.
Following Flex-DPO, the regularization methods outlined in Section 3.3, we fixed $\beta^{+}$ and systematically decreased $\beta^{-}$. As indicated by the gradient analysis in Appendix B.2.1, the gradient of rejected responses with respect to $\beta^{-}$ follows a non-monotonic trajectory, initially increasing and then decreasing. Reducing $\beta^{-}$ on the left side of this extreme point can effectively reduce the gradient magnitude. However, indiscriminately minimizing the gradient is not always advantageous. As illustrated in Figure 4, model performance does not consistently improve with an excessively small $\beta^{-}$ (see the trend with $\beta^{-}<0.08$). Over-reduction of $\beta^{-}$ risks causing the DPO algorithm to deviate from the preference learning paradigm and regress toward behavior akin to the SFT algorithm, ultimately compromising its generalization capabilities. This finding warrants further investigation, and while preliminary insights are discussed in Appendix D, deeper exploration is needed. This aspect will be addressed in future research.
Additionally, we tested other DPO variants, such as IPO and SLiC, on the MATH and SuperCLUE-Math datasets, with results presented in Table 9. Flex-DPO consistently outperforms vanilla DPO, IPO and SLiC, highlighting the effectiveness of the proposed regularization techniques.
4.4 Relative Instability of DPO Training Compared to RM Training
Figure 5: Accuracy of RM and DPO on HH-rlhf eval set over the training process.
To assess the stability and performance gap between DPO and RM training, we conducted a parallel comparison using identical datasets. Both models were built on the Baichuan2-33B architecture and trained on the HH-rlhf and UltraFeedback datasets, with evaluations conducted on the HH-rlhf and MATH datasets. The primary evaluation metric was accuracy, defined as the proportion of instances where the model correctly identified the chosen response as superior to the rejected one.
As shown in Figure 5, RM training proved to be significantly more stable, whereas DPO training exhibited notable fluctuations. These findings are consistent with the theoretical results in Section 3.4, which indicate that 3D-properties are absent in RM-based alignment methods. Furthermore, as illustrated in Figure 8 and Figure 9 in Appendix D.2, the DPO model demonstrated a higher tendency to overfit. Specifically, the sharp deceleration in accuracy improvement after the second epoch suggests that the model was overfitting the training data, highlighting the more aggressive optimization dynamics of DPO.
4.5 Suboptimality of DPO Compared to RM-based Alignment
@@
To further compare the performance of DPO and the end-to-end RM-based alignment (PPO), we tested both approaches on two datasets for poem and slogan generation. These datasets serve as ideal benchmarks for evaluating instruction-following capabilities, given their explicit and structured scoring criteria. For poem creation, the model must generate responses in accordance with specific text and tone formats based on the prompt. Evaluation metrics include five key aspects: Row Number, Words per Row, Rhythm, Tone Pattern, and Title. For slogan creation, evaluation is based on Word Count and Content. Using Baichuan2-33B for our experiments, the results, shown in Table 2, demonstrate that DPO underperforms compared to RLHF-PPO on both datasets.
5 Conclusions